work out the solutions to these simultaneous equations:
y = 2x
y = x^2 - 8

Answer by ikleyn(52778)

.
work out the solutions to these simultaneous equations:
y = 2x
y = x^2 - 8
~~~~~~~~~~~~~~~~~~~~

Left sides are identical - hence, right sides are equal

    2x = x^2 - 8,

or

    x^2 - 2x - 8 = 0,

Factor left side

    (x-4)(x+2) = 0.


One solution is  x = 4;  y = 2x = 8.       CHECK:  x^2-8 = 4^2-8 = 8.       <<<---=== correct .

Another solution is x = -2;  y = 2x = -4.  CHECK:  x^2-8 = (-2)^2-8 = -4.   <<<---=== correct .

Solved.